OpenStudy (anonymous):

Consider the equation y'' -xy' + x^2*y = 0. Find a solution y1 in the form of y = x^r, where r is a constant.

OpenStudy (jamesj):

So substitute y = x^r into the equation and you'll get an equation in r you can solve.

myininaya (myininaya):

y=Ax^r find y' and y'' and plug into the equation

OpenStudy (jamesj):

Try it. It's surprisingly easy.

OpenStudy (anonymous):

awesome thanks

myininaya (myininaya):

i don't see how it can be a solution unless i missing something

OpenStudy (anonymous):

Yeah I'm not really getting anything out of this substitution either. I don't think there are any solutions that are real.

myininaya (myininaya):

well i get \[x^r \cdot [r(r-1)x^{-2}-r+x^2]=0 \] x^r=0 when x=0 ----------- \[r(r-1)x^{-2}-r+x^2=0\] will not give us a constant r

myininaya (myininaya):

we don't want x=0 since we are looking for a solution in the form y=x^r \[r(r-1)-rx^2+x^4=0\] \[r^2-r-rx^2+x^4=0 => r^2+r(-1-x^2)+x^4=0\]

OpenStudy (jamesj):

Oh wait. Have you written down the equation correctly?

OpenStudy (zarkon):

there is probably a typo in the question

OpenStudy (jamesj):

yes -- I think so too.

OpenStudy (zarkon):

move the x^2 to y''' prob

myininaya (myininaya):

y''?

OpenStudy (zarkon):

y''

myininaya (myininaya):

lol

OpenStudy (zarkon):

too quick with the primes ;)

OpenStudy (jamesj):

yes x^2y'' - 2xy' + y = 0 would make sense.

OpenStudy (anonymous):

That's how its written, I thought it could have been a typo too but put it in case anyone saw something that I didn't. The prof is notorious for having a few typos in each assignment and not telling us till earlier on the day its due, so I'm guessing that's the case here as well.

OpenStudy (zarkon):

what you have is a nasty diff EQ

OpenStudy (jamesj):

Yes, otherwise the solution will very messy, as Zarkon just demonstrated.

OpenStudy (jamesj):

I'd assume it is the form I wrote above. Then it really IS easy!

myininaya (myininaya):

i would say there is no such solution in the form y=x^r for the dq given and then you could show there is a solution of the form to what james wrote (just in case that is what the prof meant)